Question 32

You are considering purchasing a put option on a stock with a current price of $39. The exercise price is $35, and the price of the corresponding call option is $7. According to the put-call parity theorem, if the risk-free rate of interest is 4%, and there are 60 days until expiration, the value of the put should be what?

Group of answer choices

between $2.90 and $3.00

between $2.60 and $2.70

between $2.80 and $2.90

between $2.70 and $2.80

between $2.50 and $2.60
Answer #1

Call Price + Strike Price*e^(-rt) = Put price + Stock Price
7+35*e^(-4%*60/365)=Put Price+39

Put Price =7+35*e^(-4%*1)-39 =2.77 (between $2.70 and $2.80)